OpenStudy (anonymous):

Identify the slope-intercept form of the equation of a line which contains the points (–2, –1) and (3, 4).

OpenStudy (anonymous):

\\[\frac{4-(-1)}{3-(-2)}=\frac{5}{5}=1\]slope =1

OpenStudy (anonymous):

(y - 4) = (1) (x - 3)

OpenStudy (anonymous):

\[y-(-1)=1(x-(-2)\]\[y+1=x+2\] y=x+2-1 y=x+1